# Value-Free Policy Optimization via Reward Partitioning

### v2

@@ V Qualitative Study of Prompt Completions
We conduct a qualitative analysis to illustrate the effects of response optimization strategies on real-world prompts. Specifically, we analyze a subset of prompts where both LLaMA and Mistral models demonstrate a preference for completions generated using our RPO method over baseline methods like DRO and KTO.
In each case, human preferences and reward models agree in ranking RPO higher due to its more complete, helpful, and aligned responses. DRO and KTO produce reasonable but weaker completions, often lacking nuance, informativeness, or coherence.
All examples are generated using the XL variant of each model, as it offers the highest capacity and clearest separation in quality among methods.
@@
Prompt 4: “What are some effective strategies for learning a new language quickly?”
RPO: Immersion, daily practice, spaced repetition systems, and conversational exposure are key strategies for rapid language acquisition. Leveraging media in the target language and setting achievable goals also helps maintain motivation.
DRO: Practice often, use flashcards, and talk with people.
KTO: Use apps and practice talking.
Prompt 5: “Discuss ethical concerns around AI in hiring processes.”
RPO: AI in hiring can introduce bias if trained on skewed data, raising fairness and transparency issues. Candidates may be evaluated on proxies unrelated to job performance, and lack of explainability further complicates accountability.
DRO: AI can be biased and unfair in hiring.
KTO: Bias is a concern in AI hiring tools.
Prompt 6: “What are the benefits and risks of using CRISPR for gene editing in humans?”
RPO: CRISPR offers potential cures for genetic diseases and advancements in personalized medicine. However, concerns about off-target effects, ethical boundaries, and potential misuse highlight the need for stringent oversight.
DRO: CRISPR can help cure diseases but has ethical issues.
KTO: Gene editing has risks and benefits.
Prompt 7: “How does inflation affect consumers and the broader economy?”
RPO: Inflation reduces consumers’ purchasing power, increases uncertainty, and may prompt central banks to raise interest rates. While moderate inflation can signal healthy demand, unchecked inflation erodes savings and wages.
DRO: Prices go up and people can buy less.
KTO: It makes things more expensive.
Prompt 8: “Can social media improve civic engagement? Discuss with examples.”
RPO: Social media can enhance civic engagement by facilitating access to information, enabling grassroots mobilization, and amplifying underrepresented voices. For example, platforms have played a central role in movements like the Arab Spring and Black Lives Matter.
DRO: Social media can help people stay informed and share views.
KTO: Yes, it helps people share ideas.
Prompt 9: “Compare the advantages of public versus private transportation in urban settings.”
RPO: Public transportation reduces traffic congestion and environmental impact, while private vehicles offer flexibility and convenience. Urban planning often requires balancing infrastructure investment between both modes to meet diverse needs.
DRO: Public transport is better for the environment. Cars are more flexible.
KTO: Buses are good for cities, cars are easier.
Prompt 10: “What role does sleep play in memory consolidation?”
RPO: Sleep facilitates memory consolidation by allowing the brain to process and store information from the day. REM and deep sleep stages are particularly important for transferring short-term memories into long-term storage.
DRO: Sleep helps the brain remember things better.
KTO: It helps with memory.
TABLE VII: Qualitative examples of prompt completions generated using the XL variant of each model, where both LLaMA and Mistral demonstrate preference for RPO over DRO and KTO. RPO completions are consistently more informative, well-structured, and aligned with human preferences, as reflected in UltraFeedback training.
VI Limitations
While RPO offers a simple and scalable alternative to RLHF-style optimization, it is not without limitations.
Empirical Value Estimation. RPO relies on a distributionally robust objective, where the expected value $\widehat{V}(x)$ is computed empirically over the set of candidate responses for each prompt. This estimation assumes sufficient coverage of high-reward outputs under the reference policy $\pi_{\mathrm{ref}}$. If the dataset lacks diverse or high-quality responses, $\widehat{V}(x)$ may be inaccurate, leading to suboptimal updates. In contrast to methods that explicitly train a reward or value model, RPO does not extrapolate beyond observed samples.
Dataset Quality and Coverage. The success of RPO hinges on the quality and diversity of the dataset $\mathcal{D}$. For prompts with limited variation in responses—or when rewards are noisy or sparse—the partition function $\widehat{Z}(x)$ may underestimate the reward landscape. In such settings, the algorithm may favor local modes or fail to discover beneficial deviations from the reference policy.
Dependence on $\pi_{\mathrm{ref}}$. The reference policy serves a dual role: it defines the baseline from which the updated policy diverges, and it shapes the empirical value function. If $\pi_{\mathrm{ref}}$ is poorly aligned with the target reward distribution, the optimization may be biased or less effective.
In summary, RPO’s simplicity comes at the cost of relying heavily on observed data. Addressing these limitations may involve hybridizing with reward modeling, improving data collection strategies, or introducing learned approximations of $\widehat{V}(x)$ to better generalize beyond the dataset.
VII Conclusion and Future Work
In this work, we introduced RPO—a simple, scalable, and reward-driven fine-tuning method that bypasses the need for explicit reward models or reinforcement learning loops. RPO optimizes a robust policy objective that adjusts log-probabilities to match observed preferences while staying close to a reference policy. Our method demonstrates strong empirical performance, outperforming SFT, DRO, and KTO across multiple model sizes.
RPO achieves these results while being significantly simpler to implement and train. Notably, we show that RPO benefits from higher winrates when judged by both Mistral and LLaMA LLMs, indicating strong generalization across distinct evaluators. Furthermore, ablation on the regularization parameter $\tau$ reveals robustness to hyperparameter variation, especially for larger models.
Future Work. There are several promising directions to extend this work. First, we aim to apply RPO to more powerful base models such as GPT-style architectures or frontier LLMs to assess scalability and alignment at higher capacity. Second, we plan to evaluate RPO on broader preference datasets, including real-world human feedback and multi-turn dialogues. Exploring integration with more structured exploration or dynamic reference policies may further improve coverage and robustness. Finally, hybridizing RPO with lightweight value learning or variance reduction techniques could mitigate its current dependency on dataset quality without sacrificing simplicity.

### v3

@@ V Qualitative Study of Prompt Completions
We present qualitative examples to illustrate how different optimization strategies affect response quality on real-world prompts. We focus on cases where both LLaMA and Mistral judges prefer completions generated by RPO over DRO and KTO. Across these examples, RPO produces responses that are more complete, informative, and well-structured, reflecting stronger alignment with human preferences and reward signals. In contrast, DRO and KTO generate reasonable but comparatively shallow answers, often lacking nuance or contextual depth. All examples are generated using the XL model variant, where differences between methods are most pronounced.
@@
TABLE VII: Qualitative comparison of prompt completions generated by the XL model. RPO outputs are preferred by both LLaMA and Mistral judges and are consistently more informative and coherent than those produced by DRO and KTO.
VI Conclusion
We introduced RPO, a simple and scalable fine-tuning method that normalizes observed rewards to eliminate reward models and RL loops. RPO directly supervises policies against reward-labeled completions while staying close to a reference, achieving strong performance over SFT, DRO, and KTO across model sizes. It is easy to implement, stable, and robust to hyperparameter variation, as confirmed by both Mistral and LLaMA evaluations.
While effective, RPO relies on the quality and coverage of observed data and the alignment of the reference policy, which may limit generalization beyond the dataset.
Future Work. We plan to scale RPO to larger GPT-style models, evaluate on broader datasets including human and multi-turn feedback, and explore dynamic references or lightweight value learning to improve coverage without added complexity.